OpenStudy (anonymous):

-10 sqrt 7 +12 sqrt 7

OpenStudy (anonymous):

\[10\sqrt{7}+12\sqrt{7}\]

OpenStudy (anonymous):

10+12=22 Always keep the sqrt the same. So, the answer will be 22 sqrt7.